mathslover (mathslover):

integrate (9-3x^2+4x^3 -9x^4+5x^5 )

OpenStudy (australopithecus):

split the integral up and use the basic rules

mathslover (mathslover):

\[\huge{\int ({9-3x^2+4x^3 -9x^4+5x^5})dx}\]

OpenStudy (anonymous):

integrate it one by one..

OpenStudy (australopithecus):

\[\int\limits_{}^{}(f(x) - g(x))dx = \int\limits_{}^{}f(x)dx - \int\limits_{}^{}g(x)dx\]

mathslover (mathslover):

hmn I am going to try that... wait guys

OpenStudy (lgbasallote):

just take the antiderivative. it's easier and faster

OpenStudy (australopithecus):

\[\int\limits_{}^{}x^ndx = \frac{x^{n+1}}{n+1}\]

OpenStudy (australopithecus):

\[\int\limits_{}^{}Cdx = C(x)\] where C is a constant, that is all you need to know to solve this problem

mathslover (mathslover):

\[\large{9-\frac{3x^3}{3}+\frac{4x^4}{4}-\frac{9x^5}{5}+\frac{5x^6}{6}+C}\]

OpenStudy (australopithecus):

also don't forget + C

mathslover (mathslover):

I can simplify this.. further

OpenStudy (australopithecus):

you made a mistake

mathslover (mathslover):

m i right?

OpenStudy (australopithecus):

see rule above

mathslover (mathslover):

you mean to say m wrong somewhere? @Australopithecus

OpenStudy (australopithecus):

yes Look at the rule just above your answer

mathslover (mathslover):

will there be 9x?

OpenStudy (anonymous):

when you integrate 9 its 9x

OpenStudy (australopithecus):

:L I just wanted someone to read my post haha

OpenStudy (anonymous):

and the rest are correct

mathslover (mathslover):

so it will be : \[\huge{{9x-\frac{3x^3}{3}+\frac{4x^4}{4}-\frac{9x^5}{5}+\frac{5x^6}{6}+C}}\]

OpenStudy (australopithecus):

You took that away from me Omniscience

mathslover (mathslover):

thanks a lot to both ..

OpenStudy (shane_b):

Yes...now just simplify it :)

mathslover (mathslover):

oh k so it will be: \[\huge{9x-x^3+x^4-\frac{9x^5}{5}+\frac{5x^6}{6}+C}\]

mathslover (mathslover):

thanks guys

OpenStudy (australopithecus):

You can always check your integrals with wolfram alpha but sometimes your answer might look different

OpenStudy (australopithecus):

so just be aware

OpenStudy (anonymous):

or just differentiate it to see if it matches

OpenStudy (lgbasallote):

i emvy you @mathslover you have so much people replying to your integral question...my questions are barren

OpenStudy (australopithecus):

some of your questions drive me crazy :)

OpenStudy (lgbasallote):

whuuut? but they're just algebra o.O

OpenStudy (anonymous):

the easier the question the faster it will get a respond from alot of people ;)

OpenStudy (australopithecus):

Yeah you are posting laplace transforms right? I have no idea what they are